OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches negative 8 where f of x equals x plus 9 when x is less than negative 8 and f of x equals negative seven minus x when x is greater than or equal to negative 8

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

same question you have to find the left and right limit

OpenStudy (freckles):

plug in -8 into both sides

OpenStudy (freckles):

The function that is to the left of the vertical line x=-8 is x+9. The function that is to the right of the vertical line x=-8 is -7-x.

OpenStudy (anonymous):

So would the function be 1? @freckles

OpenStudy (freckles):

you mean the limit?

OpenStudy (freckles):

-8+9=1 -7-(-8)=-7+8=1 yes both limits are 1 that means the actual limit is 1

OpenStudy (anonymous):

Ok thanks so much! @freckles

OpenStudy (freckles):

np